OpenStudy (anonymous):

Solve the word problem. A group of roommates decide to split the cost of a new sofa. If s is the cost of the sofa and n is the number of roommates, the formula determines each person’s share of the cost. If the new sofa costs $955 and 4 roommates live in the house, what is each person’s share? A. $382.00 B. $238.75 C. $221.25 D. $95.50

OpenStudy (anonymous):

the formula determins each persons *If you dont know, i didnt add the attachment, it is between "the formula" and "determins each persons"

OpenStudy (janu16):

Just divided 955 by 4

OpenStudy (anonymous):

thanks

OpenStudy (janu16):

Np

OpenStudy (anonymous):

i have more questions, can u help?

OpenStudy (janu16):

Yea if i can

OpenStudy (anonymous):

The formula P = ns gives the formula for the perimeter of a regular polygon with n sides and side length s. A regular hexagon has 6 equal sides. What is the perimeter of a regular hexagon with side length 2.8 cm?

OpenStudy (anonymous):

i would guess to multiply (2.8) 6 times?

OpenStudy (janu16):

Yess that would ne correct

OpenStudy (anonymous):

k, thanks

OpenStudy (janu16):

Np

OpenStudy (anonymous):

The formula d = rt gives the distance traveled in time t at rate r. If a bicyclist rides for 2.3 hours and travels 41 miles, what was her average speed? A. 9.43 miles per hour B. 5.6 miles per hour C. 17.83 miles per hour D. 38.7 miles per hour

OpenStudy (anonymous):

this is going to help SO many people who just google the answer. XD

OpenStudy (janu16):

If d=rt, then r=d/t 41/2.3 = 17.8261.

OpenStudy (janu16):

If ypu round it then its c

OpenStudy (anonymous):

my best guess is C. well, we got the same answer,

OpenStudy (janu16):

Yea lol

OpenStudy (janu16):

If there's more than put iton w new thread

OpenStudy (janu16):

This pne is longh

OpenStudy (anonymous):

last one, The formula d = rt gives the distance d traveled in time t at rate r. A train travels 163.2 miles in 2.4 hours, traveling at a constant rate. What is the train’s rate of travel?

OpenStudy (anonymous):

i got 71 mph

OpenStudy (janu16):

No all you have to do is multiply 163.2 and 2.4h in the formula d=rt d=163.2(2.4)= 391.68 or 68

OpenStudy (anonymous):

i would of guess's you would divide.

OpenStudy (janu16):

Yes

OpenStudy (anonymous):

for the constant rate. 163.2 / 2.4 = 71.04, round it up and its 71

OpenStudy (janu16):

I got 68 did you divide correctly?

OpenStudy (anonymous):

oh no i messed up somewhere, XD. it is 8

OpenStudy (janu16):

I got solid 68 no decimal nothing

OpenStudy (janu16):

Yea lol

OpenStudy (anonymous):

*68*

OpenStudy (anonymous):

lets see how i did...

OpenStudy (anonymous):

100% thank you so much

OpenStudy (janu16):

Np

OpenStudy (anonymous):

i guess i can try and help other people. thanks

OpenStudy (anonymous):

bye. :)

OpenStudy (janu16):

Lol yw

OpenStudy (janu16):

Byee